OpenStudy (anonymous):

2l+w=30 Solve for l.

OpenStudy (anonymous):

You have two variables and one equation. You can only express l in terms of w. Is this what you want? If so, isolate l on one side.

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

subtract w and then divide by 2

OpenStudy (anonymous):

l=(30-w)/2

OpenStudy (anonymous):

First you subtract all the terms other than l onto one side (in this case \(w\) is the term to subtract): \( 2l+w=30 \) becomes: \(2l=30-w \) Then you divide both sides by the coefficient (in this case 2 is the coefficient): \(2l=30-w \) becomes: \(l= \frac{30-w}{2}\) \(l= 15 - \frac{1}{2}w\)